# Think in Safety: Unveiling and Mitigating Safety Alignment Collapse in Multimodal Large Reasoning Model

### v1

@@ 3 Data Construction
Although existing studies have introduced several datasets for multimodal safety alignment, most of them consist of brief responses that lack explicit thought process. The supervised fine-tuning method based on these datasets struggles with retaining the inherent reasoning-chain advantages of MLRMs. Furthermore, arriving at conclusions without engaging in a safety-oriented thought process weakens the effectiveness and robustness of safety defense mechanisms. To address this issue, we propose a method for building a multimodal alignment dataset enriched with safety-oriented thought process, which offers a green channel to improve the safety of MLRMs. As a result, we aim to leverage the reasoning capabilities of MLRMs to improve the alignment with safety objectives.
As shown in Figure 5, we employ a multi-step pipeline to construct our safety dataset. After collecting various safety-related topics and image descriptions, we explicitly incorporate long CoT reasoning into the addressing process and design a filtering mechanism to finally obtain the dataset. To the best of our knowledge, it is the first safety dataset with the ability to retain reasoning chain for MLRMs.
@@
Direct 72.20 79.57 32.14 17.31 49.69
VLGuard 65.80 35.60 3.21 40.72 60.88
MIS 56.00 60.18 7.14 31.74 53.63
Ours 16.80 21.54 0.00 71.26 61.09
@@
Direct 84.80 72.26 33.57 26.95 51.26
VLGuard 72.00 25.65 10.00 75.00 51.42
MIS 75.60 58.45 11.79 79.04 51.08
Ours 20.60 11.90 0.71 97.01 54.73
Table 5: Results of supervised fine-tuning method using different datasets. The best safety scores of the MLRM on each benchmark are highlighted in bold.
@@
Due to the limitation of existing multimodal safety alignment datasets, we aim to construct a dataset enriched with safety-oriented thought process. However, considering the inherent security vulnerabilities of existing models and the development costs, it is not a practical approach to build such a dataset entirely from scratch. Therefore, we augment the existing dataset that already contains responses by adding structured and security-oriented thought process. In particular, we select two existing multimodal safety alignment datasets, VLGuard Zong et al. (2024) and SPA-VL Zhang et al. (2024), as the original data sources. Moreover, we define a set of twelve safety categories to select the original data, following the OpenAI safety guidelines, instead of using all data. More details are provided in Appendix C.1.
@@
To obtain a reliable thought process, we choose reasoning models instead of instruct models, which lack the ability to generate the thinking process with diversity and logicality. Furthermore, it is not feasible to directly utilize MLRMs to augment the thought process of the original data with their poor security performanc. Moreover, API-based proprietary models such as Claude3.7-sonnet-thinking with stronger safety mechanisms typically only return brief refusals (e.g., “I’m sorry”) when faced with some unsafe queries, offering no insight into their thought process. Therefore, we utilize the high-performing text-only reasoning model DeepSeek-R1 Guo et al. (2025) to generate safety-oriented thought process. To accommodate this model, we convert images into caption form as input in advance. In particular, we employ Qwen2.5-VL-72B Bai et al. (2025) to generate image captions to ensure that the semantic loss during modality conversion is within an acceptable range in terms of security issues, with the prompt “Please provide a detailed description of this image.
@@
To ensure the thought process generated by LRMs are truly safe, we add supplementary information related to security, inspired by the concept of deliberative alignment Guan et al. (2024). Previous methods often require models to infer implicit safety rules from large volumes of annotated examples, which suffers from low efficiency and limited generalization. Therefore, we provide clear and structured safety guidelines to directly guide the model in generating safety-oriented CoT. The guidelines not only help ensure the safety of the generated thought process, but also encourage MLRMs to think explicitly with safety considerations when producing responses. Besides, we develop the safety guidelines customized to the specific characteristics and risk profiles of each category. The image caption, question, original response, and safety guidelines are provided to DeepSeek-R1 Guo et al. (2025) to generate safety-oriented thought process. The model first analyzes the question and the caption to obtain the underlying intent, then refers to the safety guidelines to generate a safety-oriented thought process corresponding to the provided response. Prompt and examples of safety guidelines used in this step are provided in the Appendix C.2.
Thought Process Filter.
In the previous step, we obtain the original thought process. However, it is still necessary to ensure that the thought process is consistent with the corresponding response, which means that both should consistently reject or not reject the inquiry. Additionally, while the input to the model in the prior step is a caption, the input during MLRM fine-tuning is the original image. Therefore, the caption field is processed to prevent potential misinterpretation by the MLRM. The original thought process from the previous step also contain redundant elements such as headers like “Safety Chain of Thought” which need to be removed and reformatted. Given the relatively low complexity of this task, we employ DeepSeek-V3 Liu et al. (2024a) to handle this step. Specific prompts used for this step are provided in the Figure 6. After filtering, we construct the final safety datasets provided for the supervised fine-tuning of MLRMs.
@@ 4.1 Experimental Settings
To evaluate the effectiveness of our data construction, we select two existing multimodal safety alignment datasets for supervised fine-tuning as baselines that are listed as follows:.
@@
We use R1-Onevision Yang et al. (2025) and Llava-cot Xu et al. (2024) as base MLRMs for supervised fine-tuning, which clearly delineates the boundary between the thought process and the final response. We adopt supervised fine-tuning approach to the two models using the baseline datasets and our proposed dataset, respectively. The Details of the training procedure are provided in Appendix B.
@@ 4.2 Results
As shown in Table 5, the results demonstrate that the MLRMs achieve the best security performance after fine-tuning using our dataset on all benchmarks, compared with other baseline datasets. It is worth emphasizing that despite the complete absence of jailbreak attack data in our dataset, the proposed approach achieves a substantial reduction of ASR on jailbreak robustness benchmarks. Furthermore, since our dataset incorporates thought process that closely align with the data distribution used in MLRM training, the fine-tuned model preserved the original capability to generate coherent reasoning pathways. More case studies are provided in Appendix D.2. In addition, by fine-tuning the models respectively on the VLGuard, the MIS and the dataset we proposed, the safety performance of both models has been consistently improved, proving that well-designed constructing data is an effective solution to enhance the safety of the models.
@@ 5 Related Work
There are several studies that enhance the safety capabilities of MLLMs, which can be categorized into two types: the training-based method and the training-free method. The training-based method typically includes Supervised Fine-Tuning (SFT) Zong et al. (2024); Ding et al. (2025b); Li et al. (2024) and Reinforcement Learning from Human Feedback (RLHF) Zhang et al. (2024). In addition, Lee et al. (2024) introduce the weight merging approach to mitigate safety degradation. Another branch of studies incorporates additional safety components in a training-free manner. For instance, MLLM-Protector Pi et al. (2024) offers a plug-and-play detector for harmful responses, and ETA Ding et al. (2025a) proposes an inference-time alignment framework to ensure safety compliance during deployment. In contrast to these methods, our study is the first to investigate the safety behaviors of reasoning models in both jailbreak and awareness of safety scenarios.
@@ 6 Conclusion
In this work, we conduct a comprehensive evaluation and analysis of the safety of existing MLRMs, covering both jailbreak and comprehension benchmarks. The empirical results unveil several new findings, demonstrating that the safety performance of current MLRMs remains a significant concern. Furthermore, motivated by the findings, we propose a supervised fine-tuning dataset that considers explicit safety-oriented thought process. Experimental results on R1-Onevision and Llava-cot demonstrate that our dataset outperforms existing alternatives. This work represents a preliminary exploration of improving the safety of MLRMs through reasoning-based alignment. Future research will focus on developing more efficient datasets and training strategies specifically designed for MLRMs.
@@ Appendix A Evaluation Details
Our evaluation encompasses five categories of safety benchmarks, which can be broadly divided into two types: jailbreak robustness and safety-awareness. Specifically, FigStep, MMSafetyBench, and JailBreaKV fall under the jailbreak robustness benchmark, for which we report the Attack Success Rate (ASR). Meanwhile, MSSBench and SIUO are categorized as safety-awareness benchmarks, and we follow the respective safety metrics according to the original papers.
The FigStep adopts the human evaluation as safety judgment. In our experiments, we follow the classification criteria described in their paper but apply an evaluation protocol consistent with that used for the MMSafetyBench. For the JailBreaKV, we follow the same evaluation settings as their work and utilize the Llama-Guard-Vision Chi et al. (2024) as the evaluator. The benchmarks are listed as follows:
@@
$M$-SafetyBench Liu et al. (2024b) covers 13 categories and is primarily designed to evaluate the responses of MLLMs to malicious queries. In our experiments, we adopt the SD+TYPO setting, which includes a total of 1,680 samples, following the setting of Ding et al. (2025b).
@@
Appendix B Training Details
For the training procedure, we utilize the LoRA Hu et al. (2022) to fine-tune the R1-Onevision and Llava-cot based on the LLaMA-Factory [^1] https://github.com/hiyouga/LLaMA-Factory framework. All training experiments are conducted using one NVIDIA A100-80G GPUs. The hyper-parameters for supervised fine-tuning are listed in Table 6.
Hyper-Parameter Value
lora rank 8
learning rate 1.0e-5
train epoch 1.5
per_device_batchsize 1
warm up ratio 0.1
learning rate scheduler cosine
Table 6: Hyper-parameters for supervised fine-tuning.
@@ Appendix B Training Details
Qwen2.5-VL-7B(base) 93.33 11.00 52.17 93.42 6.58 50.00 51.09
R1-Onevision 86.25 15.46 50.86 86.57 10.45 48.51 49.69
@@
Appendix C Dataset Details
C.1 Data Statistics
Based on the categories provided in SPA-VL Zhang et al. (2024) and VLGuard Zong et al. (2024) datasets, we extract the raw data in the first step of construction for data preparation. In addition, we include the paired safety examples from VLGuard, where each image is matched with both a safe and an unsafe question. This design aims to improve the ability of the models to handle context-sensitive safety issues. However, the answers in SPA-VL are derived from pairwise preference data generated by large language models. Despite selecting the “chosen” responses as the final answers, a subset of them remain potential safety risks. The unsafe answers are subsequently screened and reconstructed. The statistics of data instances in each category within our dataset after the final filtering process are shown in Table 8.

### v2

@@ 3 Data Construction
Although existing studies have introduced several datasets for multimodal safety alignment, most of them Zong et al. (2024); Ding et al. (2025b) consist of brief responses that lack explicit thought process. The fine-tuning method based on these datasets struggles with retaining the inherent reasoning-chain advantages of MLRMs. Furthermore, arriving at conclusions without engaging in a safety-oriented thought process weakens the effectiveness and robustness of safety defense mechanisms. To address these issues, we propose a method for building a multimodal alignment dataset named TiS (Think in Safety), distinguished by its safety-oriented thought process and various safety-related topics. Through this method, we aim to leverage the innate reasoning capabilities of MLRM to improve its safety alignment.
As shown in Figure 5, we employ a multi-stage pipeline to construct our safety alignment dataset TiS. We begin by collecting safety-related topics and generating image captions, then explicitly incorporating long CoT reasoning into question answering. After a filtering procedure, we finally obtain the dataset. To the best of our knowledge, TiS is the first safety dataset with the ability to retain reasoning chain for MLRMs. We provide a detailed illustration of the entire pipeline in Appendix B.1.
  Jailbreak Robustness Safety-Awareness
@@
Direct 72.20 79.57 32.14 17.31 48.94
VLGuard 36.60(-35.60) 27.21(-52.36) 0.36(-31.78) 45.51(+28.20) 63.86(+14.92)
MIS 68.00(-4.20) 46.02(-33.55) 2.14(-30.00) 41.92(+24.61) 65.61(+15.67)
SPA-VL 30.00(-42.20) 35.30(-44.27) 0.36(-31.78) 42.51(+25.20) 63.60(+14.66)
TiS (Ours) 15.80(-56.40) 21.79(-57.78) 0.00(-32.14) 71.26(+53.95) 66.31(+17.37)
-w/o thought 20.80(-51.40) 16.37(-63.20) 0.00(-32.14) 63.47(+46.14) 63.92(+14.98)
@@
Direct 84.80 72.26 33.57 26.95 51.67
VLGuard 62.60(-22.20) 14.76(-57.50) 5.36(-28.21) 61.68(+34.73) 53.40(+1.73)
MIS 59.60(-25.20) 45.48(-26.78) 2.86(-3.71) 52.10(+25.15) 54.90(+3.23)
SPA-VL 41.80(-43.00) 43.63(-28.63) 0.71(-32.86) 65.27(+38.32) 57.63(+5.96)
TiS (Ours) 12.20(-72.60) 8.87(-63.39) 0.00(-33.57) 74.85(+47.90) 59.17(+7.50)
-w/o thought 28.60(-56.20) 6.67(-65.59) 2.14(-31.43) 67.66(+40.71) 53.96(+2.09)
Table 4: Results of supervised fine-tuning method using different datasets. The best safety scores of the MLRM on each benchmark are highlighted in bold and the second-best are highlighted in underline.
@@
Due to the limitation of existing multimodal safety alignment datasets, we aim to construct a dataset enriched with safety-oriented thought process. However, considering the inherent security vulnerabilities of existing models and the resource constraints, constructing such a dataset entirely from scratch is largely impractical. Therefore, we augment existing datasets with short responses by adding structured and security-oriented thought process. Specifically, we select two multimodal safety alignment datasets as our original data sources: VLGuard Zong et al. (2024) and SPA-VL Zhang et al. (2024). Rather than directly utilizing the entire datasets, we retain only the high-quality instances and ensure a balanced distribution across diverse safety-related topics.
@@
To obtain a reliable thought process, we utilize the text-only reasoning model, Deepseek-R1 Guo et al. (2025), because instruction models often struggle to produce coherent reasoning, while MLRMs tend to suffer from alignment collapse under various conditions. Meanwhile, close-source proprietary models are equipped with defensive safety guardrail, which hinders access to their internal reasoning processes particularly for safety-related questions. To address the modality gap inherent in this text-only models, the visual content is converted into detailed image captions using Qwen2.5-VL-72B Bai et al. (2025).
  R1-Onevision Safe Answer Unsafe Answer
Safe Thought 13.60% 29.40%
Unsafe Thought 14.20% 42.80%
R1-Onevision+TiS Safe Answer Unsafe Answer
Safe Thought 81.60% 15.60%
Unsafe Thought 2.60% 0.20%
LLaVA-CoT Safe Answer Unsafe Answer
Safe Thought 8.80% 30.20%
Unsafe Thought 6.40% 54.60%
LLaVA-CoT+TiS Safe Answer Unsafe Answer
Safe Thought 81.40% 11.80%
Unsafe Thought 6.40% 0.40%
Table 5: Proportion of safety in the thought process and the final answer of fine-tuned model using our dataset on FigStep.
@@
To ensure the thought process generated by LRMs is aligned with human values, we explicitly add safety guidelines in the generation process, inspired by the concept of deliberative alignment Guan et al. (2024). Specific safety standards for various domains (law, medical, etc) are also clarified. Previous methods often require models to infer implicit safety rules from large volumes of training examples, which suffers from low efficiency and limited generalization. By contrast, our method provides sufficient safety standards for safety-centric reasoning content generation, facilitating deeper safety alignment and understanding rather than the first few tokens Qi et al. (2024). More details and prompt used in this step are provided in Appendix B.3.
Thought Process Filtering.
After obtaining the safety-oriented reasoning processes, we perform an additional modification and filtering step to ensure the quality and correctness of the data, making them better suited for safety enhancement. This step addresses issues such as inaccuracies (e.g., responses that explicitly state “according to the caption”) and misalignments in reasoning. We collaboratively employ Deepseek-V3 Liu et al. (2024a) alongside human annotators to assess data quality and filter inappropriate instances.
@@ 4.1 Experimental Settings
To compare the effectiveness of our proposed dataset, we select three multimodal safety alignment datasets (MIS Ding et al. (2025b), VLGuard Zong et al. (2024), SPA-VL Zhang et al. (2024)) as our considered baselines. More details are listed in Appendix C.1.
@@
We use R1-Onevision Yang et al. (2025) and LLaVA-CoT Xu et al. (2024) as base MLRMs for safety alignment training, which demonstrates competitive performance in multimodal reasoning tasks yet falls short in safety performance. We conduct comprehensive supervised fine-tuning on both models with all above datasets and our proposed one with detailed reasoning process. The details of the training configuration are provided in Appendix C.2.
@@ 4.2 Results
As shown in Table 4, both R1-Onevision and LLaVA-CoT demonstrate improved safety alignment fine-tuned on TiS, substantially outperforming prior datasets. Specifically, our dataset TiS enhances safety performance on Figstep and SIUO by 30% compared to the best alternative baseline, effectively enabling the models to leverage their reasoning capabilities for deeper analysis and unsafe intention detection. Furthermore, since our dataset incorporates thought process that closely align with the data distribution used in MLRM training, the fine-tuned model preserved the original capability to generate coherent reasoning pathways. More case studies are provided in Appendix D.4.
4.3 Discussion
Ablation Studies.
To investigate whether the proposed thought process can further enhance the safety performance of MLRMs, we conduct experiments by removing the thought component from TiS dataset, retaining only the answer portion. As shown in Table 4, retaining the thought process leads to more substantial improvements in safety performance compared to using answer-only data, with the exception of MM-SafetyBench. For certain instances of MM-SafetyBench, the model’s deeper reasoning leads it to categorize the input as neutral, which leads to the increase of ASR.The answer-only data also exhibits favorable performance, suggesting that the construction of more comprehensive and higher-quality datasets facilitates the safe fine-tuning of MLRMs. However, this can lead to generated responses that not only lack the thought process but also consist solely of brief replies such as “I’m sorry, I can’t assist it.”.More case studies are provided in Appendix D.3. Furthermore, this also demonstrates the feasibility of leveraging the reasoning capabilities of MLRMs to enhance the alignment with safety objectives.
Analysis of Thought Process.
To further assess the impact of Tis on enhancing the security of internal reasoning, we separately evaluate the safety of both thought process and final answer, as shown in Table 5. The results demonstrate the effectiveness of TiS in safeguarding both thought & answer response, with a substantial reduction in the ASR across all categories. Notably, TiS effectively reduces the proportion of cases in which both thought process and final answer are unsafe, decreasing from 42.80% to 0.20% on R1-Onevision, and from 54.60% to 0.40% on LLaVA-CoT. Furthermore, in scenarios where unsafe content appears solely in the intermediate thinking, our fine-tuning approach also achieves a marked improvement, reducing such instances from 14.20% to 2.60%. These results indicate that TiS largely enhances the reliability and security of the model’s internal reasoning process.
@@ 5 Related Work
Current studies enhancing the safety capabilities of MLLMs can be categorized into two types: the training-based method and the training-free method. The training-based method typically includes Supervised Fine-Tuning (SFT) Zong et al. (2024); Ding et al. (2025b); Li et al. (2024) and Reinforcement Learning from Human Feedback (RLHF) Zhang et al. (2024). In addition, Lee et al. (2024) introduce the weight merging approach to mitigate safety degradation. Another branch of studies incorporates additional safety components in a training-free manner. For instance, MLLM-Protector Pi et al. (2024) offers a plug-and-play detector for harmful responses, and ETA Ding et al. (2025a) proposes an inference-time alignment framework to ensure safety compliance during deployment. In contrast to these methods, our study is the first to investigate the safety behaviors of reasoning models in both jailbreak and awareness of safety scenarios.
@@ 6 Conclusion
In this work, we systematically evaluate and analyze the safety performance of existing MLRMs, covering both jailbreak robustness and safety-awareness benchmarks. The empirical results unveil several new findings, demonstrating that the safety performance of current MLRMs remains a significant concern. Furthermore, motivated by the findings, we propose a supervised fine-tuning dataset that considers explicit safety-oriented thought process. Experimental results on R1-Onevision and LLaVA-CoT demonstrate that our dataset outperforms existing alternatives. This work represents a preliminary exploration of improving the safety of MLRMs through reasoning-based alignment. Future research will focus on developing more efficient datasets and training strategies specifically designed for MLRMs.
Limitations
Our study primarily employs MLLMs as evaluative judges due to considerations of cost-efficiency and scalability. However, relying solely on MLLMs may compromise the accuracy of safety assessments, particularly in cases involving subtle forms of unsafe content or where the model fails to correctly interpret output response. Additionally, our evaluation includes only a collection of representative MLRMs, which does not capture the full diversity and reasoning capabilities of the broader range of available MLRMs.
@@ Appendix A Evaluation Details
A.1 Benchmark and Evaluation Details
Our evaluation encompasses five categories of safety benchmarks, which can be broadly divided into two types: jailbreak robustness and safety-awareness. Specifically, FigStep, MM-SafetyBench, and JailBreaKV fall under the jailbreak robustness benchmark, for which we report the Attack Success Rate (ASR). Meanwhile, MSSBench and SIUO are categorized as safety-awareness benchmarks, and we follow the respective safety metrics according to the original papers.
For the MM-SafetyBench, FigStep and SIUO, we employ GPT-4o-mini Hurst et al. (2024) and adopt identical configurations in teir work to ensure consistency across assessments. For the JailBreaKV, we follow the same evaluation settings of their work and utilize the Llama Guard 3 Vision Chi et al. (2024) as the evaluator. However, similar to the issues observed in MIS Ding et al. (2025b), we identify shortcomings in the safe category evaluation setup of MSSBench. As a result, we employed human evaluation as an alternative. Additional details are provided in Appendix A.2. The benchmarks are listed as follows:
@@ A.1 Benchmark and Evaluation Details
MM-SafetyBench Liu et al. (2024b) covers 13 categories and is primarily designed to evaluate the responses of MLLMs to malicious queries. In our experiments, we adopt the SD+TYPO setting, which includes a total of 1,680 samples, following the setting of Ding et al. (2025b).
@@
Multimodal Large Reasoning Model Base Model
LlamaV-o1 Thawakar et al. (2025) Llama 3.2-vision-11B
LLaVA-CoT Xu et al. (2024) Llama 3.2-vision-11B
Mulberry-Llama Yao et al. (2024) Llama 3.2-vision-11B
LMM-R1 Peng et al. (2025b) Qwen2.5-VL-3B
R1-Onevision Yang et al. (2025) Qwen2.5-VL-7B
MM-Eureka Meng et al. (2025) InternVL-2.5-8B
Kimi-VL-Thinking Team et al. (2025) Kimi-VL
Skywork-R1V Peng et al. (2025a) -
QVQ-Preview Team (2024) -
Gemini2.0-Flash-Thinking DeepMind (2025) -
Claude3.7-Sonnet-Thinking Anthropic (2025) -
A.2 MSSBench Evaluation Details
In the evaluation setup of MSSBench, LLM are required to classify responses into four safety levels. For certain examples in the safe category, responses that provide helpful advice but also emphasizing potential risks are evaluated as incorrect. As shown in Figure 6, the model provides the suggestion of using a heavy-duty drill while emphasizes safety as a priority. To address this limitation, we emplyed human evaluation as an alternative.
The image, question, and response are delivered to three independent human annotators to obtain evaluation results. For the safe category, the evaluation primarily assesses whether the model exhibits unnecessary or excessive refusal. For the unsafe category, the evaluation focuses on whether the model can accurately detect and identify potential risks.
Figure 6: A failure case of MSSBench evaluation with LLM-as-Judge method
A.3 Model Details
We evaluate a total of 11 MLRMs, Table 6 summarizes the Multimodal Large Reasoning Models evaluated for safety and their corresponding base models.
Appendix B Dataset Details
B.1 Detailed Pipeline
Step 1: Data Preparation.
Based on the categories provided in SPA-VL Zhang et al. (2024) and VLGuard Zong et al. (2024) datasets, we extract the raw data in the first step of construction for data preparation. In addition, we include the paired safety examples from VLGuard, where each image is matched with both a safe and an unsafe question. This design aims to improve the ability of the models to handle context-sensitive safety issues. However, the answers in SPA-VL are derived from pairwise preference data generated by large language models. Despite selecting the “chosen” responses as the final answers, a subset of them remain potential safety risks. The unsafe answers are subsequently screened and reconstructed.
Step 2: Image Description Generation.
To accommodate the input requirements of DeepSeek-R1 Guo et al. (2025), we convert images into caption form as input in advance.In particular, we employ Qwen2.5-VL-72B Bai et al. (2025) to generate image captions to ensure that the semantic loss during modality conversion is within an acceptable range in terms of security issues, with the prompt “Please provide a detailed description of this image.”.
Step 3: Safety Thought Process Generation.
After obtaining the image captions in the second step, the image caption, question, original response, and safety guidelines are provided to DeepSeek-R1 Guo et al. (2025) to generate safety-oriented thought process. The model first analyzes the question and the caption to obtain the underlying intent, then refers to the safety guidelines to generate a safety-oriented thought process corresponding to the provided response.
Step 4: Thought Process Filter.
The primary objective at this step is to filter and transform the original thought process. In particular, the caption field is processed to prevent potential misinterpretation by the MLRM. The original thought process from the previous step also contain redundant elements such as headers like “Safety Chain of Thought” which need to be removed and reformatted. Specific prompts used for this step are provided in the Figure 7.
@@ B.1 Detailed Pipeline
Qwen2.5-VL-7B(base) 93.33 10.67 52.00 93.42 2.63 48.03 50.02
R1-Onevision 86.60 15.12 50.86 86.57 7.46 47.01 48.94
@@
B.2 Dataset Statics
The statistics of data instances in each category within our dataset after the final filtering process are shown in Table 8.